Find the area to the left of critical value = 2.500 when the sample size is 29. [Hint: One-tailed test]


1
Expert's answer
2021-06-15T06:53:46-0400

Given sample size = 29, degrees of freedom = n-1= 29-1=28 Looking at t table,


The critical value:

"t_{.99}=2.500"

So, the area to the left of critical value=99%